Problem 4-10
Present and Future Values of Single Cash Flows for Different Interest Rates

Use both the TVM equations and a financial calculator to find the following values. Round your answers to the nearest cent. (Hint: Using a financial calculator, you can enter the known values and then press the appropriate key to find the unknown variable. Then, without clearing the TVM register, you can "override" the variable that changes by simply entering a new value for it and then pressing the key for the unknown variable to obtain the second answer. This procedure can be used in parts b and d, and in many other situations, to see how changes in input variables affect the output variable.)

  1. An initial $500 compounded for 10 years at 9.2 percent.
    $   
  2. An initial $500 compounded for 10 years at 18.4 percent.
    $   
  3. The present value of $500 due in 10 years at a 9.2 percent discount rate.
    $   
  4. The present value of $500 due in 10 years at a 18.4 percent discount rate.
    $  
Answer #1

a. Information provided:

Present value= $500

Time= 10 years

Interest rate= 9.2%

The future value is calculated by entering the below in a financial calculator:

PV= -500

N= 10

I/Y= 9.2

Press the CPT key and FV to compute the future value.

The value obtained is 1,205.58.

Therefore, the future value is $1,205.58.

b. Information provided:

Present value= $500

Time= 10 years

Interest rate= 18.4%

The future value is calculated by entering the below in a financial calculator:

PV= -500

N= 10

I/Y= 18.4

Press the CPT key and FV to compute the future value.

The value obtained is 2,706.99

Therefore, the future value is $2,707.

c.Future value= $500

Time= 10 years

Interest rate= 9.2%

The present value is calculated by entering the below in a financial calculator:

FV= -500

N= 10

I/Y= 9.2

Press the CPT key and PV to compute the present value.

The value obtained is 207.37.

Therefore, the present value is $207.37.

d.Information provided:

Future value= $500

Time= 10 years

Interest rate= 18.4%

The present value is calculated by entering the below in a financial calculator:

FV= -500

N= 10

I/Y= 18.4

Press the CPT key and FV to compute the present value.

The value obtained is 92.35.

Therefore, the present value is $92.35.

In case of any query, kindly comment on the solution.